f³f¹(x) = √secx.x³
1
Expert's answer
2020-06-01T18:22:33-0400
"f(x)=\sqrt{\sec x}\cdot x^3"

"{df\over dx}={x^3\over 2\sqrt{\sec x}}(-{-\sin x\over \cos^2x})+3x^2\sqrt{\sec x}="

"=x^2{x\tan x+6\over 2\cos x\sqrt{\sec x}}=x^2{x\tan x+6\over 2\sqrt{\cos x}}="

"=x^2\sqrt{\sec x}\cdot{x\tan x+6\over 2}"